OpenStudy (anonymous):

Simplify the expression: √72a^5

OpenStudy (anonymous):

This can be rewritten as\[\sqrt {72a^5} = \sqrt{36 * 2} \sqrt {a^5} = 6\sqrt{2} \sqrt{a^4 a} = 6a^2 \sqrt{2a}\]